OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment that joins points (5,2) and (-3,8)

OpenStudy (carson889):

Midpoints = \[\frac{ x _{2}+x _{1} }{ 2 },\frac{ y _{2}+y _{1} }{ 2 }\] \[\left( \frac{ -3+5 }{ 2 } ,\frac{ 8+2 }{ 2 }\right)=\left( 1,5 \right)\]